A force of 300 N acts on a 2.5 m long and 10 kg rod initially at rest. what is the final angular momentum of the rod if the force acts on it for 8 seconds?
1
Expert's answer
2020-12-18T11:10:51-0500

"solution"

"if F=300N"

"r=2.5m"

"m=10kg"

"u=0m\/s"

"t=8s"

"from F=ma"

"300=10*a"

"a=300\/10"

"a=30m\/s^2"

"from v=u+at"

"v=0+30*8"

"v=240m\/s"

"linear momentum" "p = m*v."

"=10*240"

"=2400kgm\/s"

the final angular momentum "L" "=" distance of the object from a rotation axis"(r)" "*" the linear momentum"(p)"

"L=r*p"

"=2.5m*2400kgm^2\/s"

"=6000kgm^2\/s"